### v1

@@ 4.3 Discussion on $r$
Figure 6 demonstrates the performance with different $r$. The outcome reveals that for dynamic datasets, the best rank option falls around $\frac{T}{2}$ while for static datasets, the best value of $r$ (also known as connecting factor) is very small. Dynamic datasets exhibit significant variations in content across different time steps. Conversely, static datasets involve the duplication of frames, where each frame is replicated $T$ times as input to the network. Consequently, this duplication results in the attention map being subjected to a low-rank condition. Considering a comprehensive assessment of both experimental outcomes and theoretical insights, our conclusion asserts that the precision of the attention map in faithfully mirroring the input tensor is unnecessary. Such fidelity does not inherently facilitate the effective concentration on pivotal components. Furthermore, our findings unveil an intriguing phenomenon: even in scenarios where SNNs exhibit considerable sparsity, the rank of the attention map can still remain markedly low. This revelation implies the presence of untapped optimization potential within the realm of SNN sparsity.
@@ 4.4 Ablation Study
In this section, we conduct ablation experiments from both global and local perspectives to verify the reasonableness of PFA. We first try adding different numbers of PFA layers to the network to see how PFA improves classification accuracy. Then we set the three projections obtained inside the PFA to a constant matrices with each element assigned to 1 respectively to eliminate the attention on this dimension to explore the impact of different dimensions of attention on the classification results.
We insert different numbers of PFAs into the VGG network and obtain the results in Figure 7 (left). In this experiment, $r$ is fixed at 8. We can observe that as more and more PFA layers are inserted, the accuracy of the network is getting higher and higher. Even inserting only one layer can improve the accuracy of the network.
In order to explore how attention in temporal-channel-spatial dimensions affects classification accuracy, we conduct an ablation by fixing corresponding projection matrix obtained by LPST to a all-one tensor. As shown in Figure 7 (right), whichever dimension is ablated will adversely affect the results. Ablating temproal and ablating channel have similar effects. This means that PFA treats temporal and channel dimensions equally. However, when we ablate the spatial dimension, the accuracy of the network drops significantly (79.60%) but still outperforms the baseline (77.90%). Whichever dimension we ablate, the performance drops. This observation highlights that PFA effectively directs attention to distinct regions across multiple dimensions, aligning with the intended purpose of attention modules. Specifically, PFA exhibits attention distribution across all three dimensions, albeit with varying weights. Notably, the spatial dimension emerges as the most significant, followed by the time and channel dimensions, in terms of attention importance.
Figure 7: Results of ablation experiments. The left one is ablation study on PFA layers. The x-axis represents the amount of PFA inserted. The right one is the result of ablation study on temporal-channel-spatial ablation. Full denotes the original PFA without any modification as a reference. w.o. T/C/S denotes setting the temporal/channel/spatial projection to a constant matrix with each element assigned to 1.
@@ 4.5 Attention Visualization
In order to visually represent the attention distribution of the PFA module, we have generated an attention map in PFA and depicted it in Figure 8. The attention maps corresponding to the temporal and channel dimensions in the figure are obtained from the deep layer of the network. This choice was made because deep layers tend to provide a more accurate reflection of the attention’s role. Conversely, the spatial attention map is obtained from the input layer of the network, where the shallow graphic features are more distinct and conspicuous, facilitating analysis and observation. Figure 8 also illustrates the distinctive attention distributions in the temporal, channel, and spatial dimensions achieved by PFA. This finding aligns with the conclusions drawn from our ablation study, further validating our earlier assertions.
@@ 5 Conclusions
In this paper, we introduce the novel approach named PFA, which combines learning and generation of a temporal-channel-spatial attention map through CP decomposition. Specifically, we propose the LPST to learn a projection of the input tensor that can extract essential information for AMC to compose the attention map. In the AMC, we utilize the reverse process of CP decomposition to generate the attention map. After analyzing the rank of the raw input tensor and the best option for connecting factor $r$, we propose a principle to search for the best $r$. Remarkably, our findings suggest that even a low rank can lead to outstanding performance. Regardless the form of the dataset, dynamic or static, PFA achieves the SOTA performance. Additional ablation studies prove the reasonableness of PFA while visualization of PFA attention distribution helps us intuitively understand the significance effect of PFA.
Figure 8: Attention visualization on CIFAR10DVS dataset. Temporal and channel parts are taken from the deep layer of VGGSNN. The spatial layer visualizations are extracted from the shallow layer of the network due to the inherent complexity of the information contained in the deeper layers, which may impede human comprehension and hinder the interpretability of our analysis.

### v2

@@ 4.4 Discussion on $R$
Figure 8 demonstrates the performance with different $R$. The outcome reveals that for dynamic datasets, the best rank option falls around $\frac{T}{2}$ while for static datasets, the best value of $R$ (also known as connecting factor) is very small.
Dynamic datasets exhibit significant variations in content across different time steps. Conversely, static datasets involve the duplication of frames, where each frame is replicated $T$ times as input to the network. Consequently, this duplication results in the attention map being subjected to a low-rank condition. Considering a comprehensive assessment of both experimental outcomes and theoretical insights, our conclusion asserts that the precision of the attention map in faithfully mirroring the input tensor is unnecessary. Such fidelity does not inherently facilitate the effective concentration on pivotal components. Furthermore, our findings unveil an intriguing phenomenon: even in scenarios where SNNs exhibit considerable sparsity, the rank of the attention map can remain markedly low. This revelation implies the presence of untapped optimization potential within the realm of SNN sparsity.
It is worth noting that on the CIFAR10-DVS dataset, when $R=1$, the performance of PFA is comparable to TCJA (see Table 2 and does not show significant advantages. This phenomenon strongly confirms our previous statement that TCJA is a "rank-1" attention, a specal case of PFA, highlighting the necessity and advantages of non-"rank-1" attention modules.
@@ 4.5 Ablation Study
Figure 9: Ablation study on PFA layers. The x-axis represents the amount of PFA inserted. The accuracy improves with the insertion of PFA layer.
Figure 10: Ablation study on attention dimensions. T/C/S denotes temporal/channel/spatial correspondingly. SC denotes spatial-channel joint attention and so on. Full means paying attention to three dimensions at the same time, i.e. a standard PFA module.
Figure 11: Attention visualization on CIFAR10DVS dataset. Temporal and channel parts are taken from the deep layer of the VGG network. The spatial layer visualizations are extracted from the shallow layer of the network due to the inherent complexity of the information contained in the deeper layers, which may impede human comprehension and hinder the interpretability of our analysis.
In this section, we perform ablation experiments from both global and local perspectives to validate the efficacy of PFA on CIFAR10DVS [29] Initially, we vary the number of added PFA layers in the network to assess the impact on classification accuracy. Subsequently, we set the three projections obtained within the PFA to constant matrices, with each element assigned a value of 1, aiming to eliminate attention to specific dimensions. This helps explore the influence of different attention dimensions on classification results. Finally, for a more comprehensive insight into the influence of PFA on the training process, we conduct ablation experiments specifically involving the insertion of PFA.
We incorporate varying numbers of PFAs into the VGG network and present the results in Figure 9. In this experiment, the value of $R$ is held constant at 8. It is evident that the network’s accuracy exhibits a noticeable improvement with more PFA layers. Remarkably, even the addition of a single layer leads to enhanced accuracy in the network.
In order to explore how attention in temporal-channel-spatial dimensions affects classification accuracy, we conduct an ablation by fixing the corresponding projection matrix obtained by LPST to an all-one tensor. As shown in Figure 10, whichever dimension is ablated will adversely affect the results. Ablating temporal and ablating channels have similar effects. This means that PFA treats temporal and channel dimensions equally. However, when we ablate the spatial dimension, the accuracy of the network drops significantly (79.60%) but still outperforms the baseline (77.90%). Whichever dimension we ablate, the performance drops. When we restrict the focus dimensions to only one, namely, spatial, temporal, or channel, we observe another decrease in the network’s performance compared to when attention is distributed across two dimensions. The retention of spatial attention significantly impacts accuracy, followed by temporal attention. This observation highlights that PFA effectively directs attention to distinct regions across multiple dimensions, aligning with the intended purpose of attention modules. Specifically, PFA exhibits attention distribution across all three dimensions, albeit with varying weights. Notably, the spatial dimension emerges as the most significant, followed by the time and channel dimensions, in terms of attention importance.
@@ 4.6 Attention Visualization
In order to visually represent the attention distribution of the PFA module, we have generated an attention map in PFA and depicted it in Figure 11. The attention maps corresponding to the temporal and channel dimensions in the figure are obtained from the deep layer of the network. This choice was made because deep layers tend to provide a more accurate reflection of the attention’s role. Conversely, the spatial attention map is obtained from the network’s input layer, where the shallow graphic features are more distinct and conspicuous, facilitating analysis and observation. Figure 11 also illustrates the distinctive attention distributions in the temporal, channel, and spatial dimensions achieved by PFA. This finding aligns with the conclusions drawn from our ablation study, further validating our earlier assertions.
@@ 5 Conclusions
In this paper, we introduce a novel attention module named PFA, which combines learning and the generation of a temporal-channel-spatial attention map through CP decomposition. Specifically, we propose the LPST to learn a projection of the input tensor that can extract essential information for AMC to compose the attention map. In the AMC, we utilize the reverse process of CP decomposition to generate the attention map. After analyzing the rank of the raw input tensor and the best option for connecting factor $R$, we propose a principle to search for the best $R$. Remarkably, our findings suggest that even a low rank can lead to outstanding performance. On classification tasks, regardless of the form of the dataset, dynamic or static, PFA achieves the SOTA performance. When we extend our evaluation of PFA to image generation tasks, it still shows its capacity to enhance performance on this task as well. Furthermore, additional ablation studies bolster the rationale behind PFA’s efficacy, while visualizing the attention distribution of PFA offers intuitive insights into its substantial impact.
In the future, we plan to extend the applicability of the PFA model to various demanding intelligent applications and deploy the model on a neuromorphic chip.